the sum of the number of A's and B's on a math test is 15. There are more B's than A's, and the sum of the squares of the two numbers is 113. Find the number of A's and number of B's.
Answer by stanbon(75887)

the sum of the number of A's and B's on a math test is 15. There are more B's than A's, and the sum of the squares of the two numbers is 113. Find the number of A's and number of B's.
-----
A + B = 15
B > A
A^2 + B^2 = 113
----
Find 2 perfect squares whose sum = 113 and whose difference = 15
64 + 49 = 113
64 - 49 = 15
-----
Since B > A, B = 8 A = 7
------------
Cheers,
Stan H.